Find all values of z, e^(2iz)=-1
1
Expert's answer
2011-11-03T13:22:49-0400
-1=e^i(pi+2*pi*k) where k - integer
so we have e^(2iz)=-1=e^i(pi+2*pi*k)
Hence 2iz=i(pi+2*pi*k)
=> z=(pi+2*pi*k)/2& where k -integer
pi=3.14.15926...